OpenStudy (erinweeks):

Solve the system of equations using matrices. Use Gaussian elimination with back-substitution.

OpenStudy (erinweeks):

x + y + z = -5 x - y + 3z = -1 4x + y + z = -2 A. {( 1, -4, -2)} B. {( -2, 1, -4)} C. {( 1, -2, -4)} D. {( -2, -4, 1)}

OpenStudy (john_es):

(A) (1,-4,-2). Do you need the solution step by step?

OpenStudy (erinweeks):

No ..!